OpenStudy (anonymous):

In 1990, people in the United States spent about $285.7 billion on recreation. In 1995, they spent $402.5 billion. Using a linear equation in slope-intercept form to model the amount (in billions of dollars) spent on recreation since 1990, predict the amount of money spent on recreation in the United States in 2000. SOME ONE HELP PLEASE?

OpenStudy (anonymous):

y = 23.36 x + 285.7 In the year 2,000: y = 23.36 * 10 + 285.7 = $ 521.1 billions